OpenStudy (ashy98):

4 subtracted from x is equal to x divided by 3. What is x?

OpenStudy (ranga):

x - 4 = x/3 multiply both sides by 3 3x - 12 = x add -x to both sides 2x - 12 = 0 add 12 to both sides 2x = 12 divide both sides by 2 x = 6

OpenStudy (anonymous):

12-3x=x 4x=12 x=3

OpenStudy (ranga):

The equation looks like this, correct? \[\Large x - 4 = \frac{ x }{ 3 }\]

OpenStudy (ashy98):

yes it does!(:

OpenStudy (ranga):

Then x = 6 is the correct answer. Put x = 6 on the left and you get 6 - 4 = 2 Put x = 6 on the right and you get 6/3 = 2 LHS = RHS and therefore x = 6 is the correct answer.

OpenStudy (ashy98):

okay thank you !(:

OpenStudy (ranga):

you are welcome.